OpenStudy (anonymous):

what is the slope of the line that goes through (0,0) and (5,-1)?

hartnn (hartnn):

slope of a line through origin is just m = y/x here, y=-1, x=5 just plug in! :)

OpenStudy (anonymous):

\[(-1-0) / (5-0)\]

OpenStudy (anonymous):

-1/5 = slope.

OpenStudy (anonymous):

:)